determine the constant that should be added to the binomial so that it becomes a perfect square trinomial. factor the trinomial. x^-5x
Found 2 solutions by stanbon, fractalier:
Answer by stanbon(48558)

determine the constant that should be added to the binomial so that it becomes a perfect square trinomial. factor the trinomial.
x^2-5x+(5/2)^2
Factors are (x-(5/2))(x-5/2))
Cheers,
Stan H.

Answer by fractalier(2101)

You probably meant
x^2 - 5x
The number we add is
half of -5, squared...that is
[(1/2)(-5)]^2 = (-5/2)^2 = 25/4
and is always positive.
So from x^2 - 5x + 25/4, we factor it as
(x - 5/2)^2.
Notice the -5/2 is half of the original -5 from above...